Which of the following functions is uniformly continuous on the given set?

a) f(x)= x3 on [1,infinity)

b) f(x)= x3 on [1,5]

c) f(x)=1/x on (0,1)

d) f(x)=1/x on [0,1]

e) All of the above


1
Expert's answer
2021-02-08T16:56:45-0500

Which of the following functions is uniformly continuous on the given set?

a) f(x)= xon [1,infinity)

b) f(x)= xon [1,5]

c) f(x)=1/x on (0,1)

d) f(x)=1/x on [0,1]

e) All of the above


A function is said to uniformly continuous a set say, S, for every "\epsilon>0" , "\exists" "\delta>0" such that "x,a\isin S" and "\mid x-a\mid<\delta" then "\mid f(x)-f(a)\mid <\epsilon"


Also a function is said to be uniformly continuous A function is said to uniformly continuous a set say, S, for every "\epsilon>0" , "\exists" "\delta>0" such that "x,a\isin S" and "\mid x-a\mid<\delta" then "\mid f(x)-f(a)\mid \ \ge \epsilon"





(c) d) f(x)=1/x on (0,1)

suppose f is uniformly continuous

then pick "\epsilon=1" then "\exists" "\delta>0: \forall" "x,a \isin (0,1)" with "\mid x-a\mid<\delta"

the "\mid {1 \over x} - {1 \over a}\mid<\epsilon"

Pick x(0,1) with "x<\delta" then set "a = {x \over 2}"

"\mid x-a\mid< \mid x- {x \over 2}\mid=\mid {x \over2}\mid< {\delta \over 2}"

"\mid {1 \over x} - {1 \over a}\mid=\mid {1 \over x} - {1 \over{x \over 2}}\mid=\mid {1 \over x} - {2 \over x}\mid=\mid -{1 \over x} \mid={1 \over x}"


Since x is in the interval (0,1) it implies that "{1 \over x}>1"

therefore the function is not uniformly continuous


d) f(x)=1/x on [0,1]

this has the same prove as (c) it is also not not uniformly continuous


b) f(x)= xon [1,5]

"\epsilon >0" then Let "\delta={\epsilon \over 108}"


Supppose "\mid x-a\mid< \delta" then "\mid f(x) -f(a)\mid =\mid x^3 - a^3\mid = \mid x-a\mid\mid x^2+xa+a^2\mid<\epsilon"

"<{ \epsilon \over 108}.108 = \epsilon"


"\therefore" the function uniformly continuous




a) f(x)= xon [1,infinity)